# UNK-VQA: A Dataset and A Probe into Multi-modal Large Models’ Abstention Ability

### v3

@@ I Introduction
Figure 1: Left: Five perturbation types from our UNK-VQA dataset and their corresponding exemplars. Right: The unanswerability ratio with respect to each perturbation type.
@@
Despite the crucial function of these benchmarks during the early phase of exploration, recent progress has largely been driven by the emergence of large models [7, 8, 9]. These large models overshadow the common training-and-evaluation paradigm in various domains, including VQA. For example, GPT-4 [8] achieves a zero-shot accuracy of $\sim$80on the VQA v2 dataset [1], significantly outperforming most supervised approaches. Pushing the limits on these traditional well-studied benchmarks thus appears to yield marginal contributions. While effective, these large foundation models, however, are notorious for being trustworthy. One demonstration is that they are less capable of abstaining from answering questions that cannot be answered or are beyond their scope of knowledge [10].
@@
This paper proposes a dataset that teaches machines to identify and refuse unanswerable visual questions. Arguably, annotating a large-scale dataset with unanswerable questions is extremely laborious and difficult to control. To address this challenge, we suggest introducing perturbations to existing data and curating hard instances that can potentially deceive models. Based on the most popular VQA v2 dataset [1], we apply five different types of nuanced perturbations to either the given image or the question, as shown in Fig. 1. To assure high quality, we engage the Amazon Mechanical Turk (MTurk) workers to label them manually. In short, the total number of participating workers is $>$4,000, and each instance is annotated by a minimum of three workers. This finally amounts to 10K instances in our dataset, which we name as UNK-VQA. As can be observed from Fig. 1, masking the most relevant object (I-2) often leads to more difficulties in performing question answering. Unlike existing approaches [11, 12, 13] that employ unsupervised learning techniques to address unanswerable visual questions, our proposed dataset enables model training towards more generalizable VQA methods.
@@
We evaluate the performance of multiple multi-modal large models including Otter [14], Open-Flamingo [9], and InstructBLIP [15]. Our findings indicate that: 1) In terms of zero-shot performance, our UNK-VQA dataset reveals limitations in these large models, unlike the impressive results observed in previous general VQA datasets. 2) For few-shot performance, when we introduce more instances to these models, consistent improvements in their performance can be observed. Additionally, we also propose a simple method to enhance VQA models with the ability of abstention. The key to our method lies in the design of selection functions, which we have developed in several variants, such as a binary classifier and entropy-based methods. Considering the resource-intensive nature of fine-tuning large models, we apply this method to several conventional and recent medium-sized VL Transformer models.
@@
This dataset reflects the fact that existing multi-modal large models, in comparison to their counterparts in the language domain, are not as omnipotent. It sheds light on various issues beyond the challenge of abstaining from answering unanswerable questions. Problems such as consistency in VQA [16], factuality, and hallucination [8] further emphasize the need for further research in this domain.
@@
  • We introduce a straightforward method for training VQA models that enables them to handle unanswerable questions.
Figure 2: Illustration of two text-based perturbations. We utilize a strong multi-modal model, i.e., BLIP [17] to generate answers (Base-A) to the perturbed question (Pert-Q). In the case of word replacement examples, when we replace the anchor noun word with an alternative, the baseline still generates a reasonable answer, even though the modified question becomes unanswerable. Regarding the semantic negation examples, each question can have an uncoutable number of potential answers.
@@ II-A Perturbation Procedure and Quality Control
In this paper, we introduce five distinct types of perturbations that are applied to either the given question or image. We provide a detailed explanation of each of these perturbations below.
@@
Figure 4: Sunburst distribution of the first four words in the UNK-VQA questions. Most questions begin with the word ‘what’.
@@ II-B Human Labeling
Comparison with related datasets. We compare our UNK-VQA dataset with several representative and related datasets, and the results are summarized in Table I. It is evident that only RGQA and our UNK-VQA introduce perturbations to VQA instances. However, UNK-VQA has two distinct advantages over RGQA: 1) We employ a diverse crowd-sourcing labeling approach, involving over 4,000 annotators, while RGQA only utilized 8 annotators. 2) Unlike RGQA, our dataset allows for training on unanswerable VQA pairs, whereas RGQA is designed solely for testing purposes. Furthermore, it is worth noting that the images in the VizWiz dataset were captured by blind individuals. Consequently, many images with unanswerable questions exhibit a significant semantic gap between the questions and the visual content, which can result in shortcut predictions by models. In contrast, our dataset ensures a strong semantic coherence between questions and images for all instances. To facilitate training or fine-tuning on UNK-VQA, we further divided the dataset into training, validation, and testing sets, consisting of 7K, 1K, and 2K instances, respectively.
@@ II-C Data Analysis
Figure 5: Labeling confidence and consensus among three annotators (y-axis: counted numbers). The majority of annotators exhibit a higher level of confidence and are able to reach a strong consensus regarding the answerability of the questions.
@@
Figure 6: Answer and Reason interplay of unanswerable questions (left) and answer shift of answerable questions (right). The left side illustrates the interplay between answers and reasons for unanswerable questions, denoted as A* and R* respectively (as defined in Sec. II-B). The most frequent reasons observed are Being unclear to comprehend, while the most common answer is I don’t know. On the right side, the figure depicts the answer shift of answerable questions. Notably, many answers shift from the Original ground-truth answers to the Baseline predicted answers.
@@ III Robustness of Large Models towards Unanswerable Questions
Given the remarkable performance of large models across various tasks, we aim to investigate their effectiveness on our UNK-VQA. Initially, we opt to explore the capabilities of the widely used ChatGPT model [^2] https://openai.com/chatgpt., as the multi-modal capabilities of GPT-4 [8] have not yet been publicly disclosed. To adapt the dataset to ChatGPT’s text-only requirement, we employ the strong image caption model BLIP-2 [7] to generate detailed captions for each image. These captions are then used as the supporting document when asking questions to ChatGPT. However, the outcomes are far from convincing. The generated captions failed to accurately depict the image content, often missing crucial objects or neglecting nouns mentioned in the question. Consequently, ChatGPT deemed most questions as ‘unanswerable’, leading to unexpected results. In light of these limitations, we present to directly employ multi-modal large models for our evaluation.
III-A Evaluation Setup
@@
  • Open-Flamingo-MPT-7B follows similar architecture with Flamingo [9] but incorporates open-sourced vision and language encoder components, specifically CLIP [20] and MPT [27]. The model is trained using a combination of LAION-2B [28] and Multimodal C4 [29], with the weights of both the vision and language encoders frozen. To align multimodal information, a trainable Perceiver resampler [30] is used to extract patch features. These patch features are then inputted into the dense cross-attention modules to connect the vision features to the corresponding text tokens.
@@
  • Instruct-BLIP-Vicuna-7B [15] is built on the pre-trained BLIP-2 [7]. It leverages a dataset in an instruction-tuning format, which has been transformed from publicly available datasets.
@@
TABLE II: Zero-shot performance of three large multi-modal models on three distinct settings.
@@
  • Open-Ended with Hint (OEH) involves an extra hint, namely Note, the question may be unanswerable., compared to OE.
@@
Options:
A. #option A
B. #option B
C. #option C
D. #option D
@@
#question Note, it is possible that the question is unanswerable based on the image.
Note that in the $k$-shot setting, the above prompts are repeated $k$ times combined with different pre-defined examples, which serve as few-shot prompts. Additionally, the entire inference process described above is performed on 4 NVIDIA A40 GPUs with distributed data parallel.
TABLE III: OoS (Out-of-Scope) Response Ratio of Otter-MPT-7B. Generally, we observe that when the accuracy (Acc) of the model is higher, the OoS response ratio tends to be lower.
@@
54.83${}_{{\color[rgb]{0,0,1}+9.57}}$ 54.64 54.82${}_{{\color[rgb]{0,0,1}+0.25}}$
III-B Evaluation Results
Zero-shot Performance We present the zero-shot results in Table II and observe that Instruct-BLIP-Vicuna-13B outperforms other models in both BY and OE settings. Notably, the performance improvements of Instruct-BLIP-Vicuna-13B compared to Instruct-BLIP-Vicuna-7B highlight the advantages of scaling up the LLM backbone. Another observation is that Open-Flamingo exhibits superior performance compared to the other three models in the MC setting. One possible explanation for this is that, in contrast to BY and OE, MC deviates more from the objective of instruction tuning, e.g., open-domain VQA and image captioning. Consequently, Open-Flamingo receives less training in instruction tuning, which sets it apart from the other models.
@@ III-B Evaluation Results
Few-shot Performance We show the results of this experiment in Fig. 7. One can see that the model performance generally improves consistently across most experiments. However, one special case with Otter is that the increased number of shots leads to a decline in results. Upon examining the test cases of Otter, we discovered that having more prompted examples often leads to an increase in out-of-scope responses. For example, some models tend to respond I do not know or attempt to explain the unusual nature of the image. These are instances where there is no explicit choice of being answerable or unanswerable in the responses. Additionally, the model’s performance significantly varies based on the number of answerable and unanswerable examples, as illustrated in Table III.
Effectiveness of Post-hint We further investigate whether the post-hint explanations assist in determining the answerability of visual questions. Based on this idea, we provide hints for the unanswerable instances and present the final results in Table IV. The table demonstrates that, in most cases, these post-hints aid the model in comprehending the query more effectively, especially for examples with fewer shots.
@@ IV-A Preliminary
where $\hat{g}(x)$ indicates how confident the classifier $f$ is on the input $x$, and $\theta$ controls the overall prediction versus abstention level.
@@ IV-C Variants of Selective Functions
Classifier-based approach involves the introduction of an additional binary classifier, which facilitates the determination of question answerability by comparing the predicted score with a predefined threshold $\theta$. Specifically, we re-utilize the fused multi-modal feature $\mathbf{x}$ as inputs to this classifier, represented as $\sigma(\mathbf{W}_{b}\mathbf{x}+\mathbf{b})$, where $\sigma$ represents the sigmoid function.
Entropy-based approach compares the entropy of the predicted logits with the threshold $\theta$. If the entropy of the predicted logits, denoted as $H(\mathbf{\pi})=-\sum_{j}^{|\mathcal{A}|}\pi_{j}\log(\pi_{j})$ [^4] In our implementation, we ease the constraint of the infinite range of entropy by comparing the maximum logit with the specified threshold., exceeds $\theta$, indicating a higher degree of randomness, the question is classified as unanswerable. During training, we assign a uniform distribution, $\{\frac{1}{|\mathcal{A}|},\frac{1}{|\mathcal{A}|},...\}$, to label the unanswerable questions, thereby reducing the model’s confidence to answer such questions. This approach aims to effectively handle cases where the model should refrain from providing a definitive answer.
TABLE V: Model performance on UNK-VQA. CLS and ENT represent classifier- and entropy-based approaches in Sec. IV-C, respectively. We did not implement the entropy-based variant of BLIP because it would have required more computational power due to the generative decoding strategy. Instead, we chose to include an additional label, unanswerable, as a generative candidate during both training and inference.
Model CLS ENT Validation Test
[…]
@@
V Experiments
V-A Experimental Settings
Below, we provide details on the evaluation metrics and baselines that were employed in our experiments.
Metrics For our evaluation, we adopted three metrics:
[…]
    $\begin{cases}P_{j}=\frac{TP_{j}}{TP_{j}+FP_{j}},\\ R_{j}=\frac{TP_{j}}{TP_{j}+FN_{j}}.\end{cases}$
    The final metric takes the class imbalance into consideration,
    where $\psi_{j}$ denotes the list of ground-truth samples with answer $j$.
Baselines We selected three popular baselines to evaluate the effectiveness of the proposed method.
  • LXMERT [36] is built upon the Transformer encoders to learn the connections between vision and language. It undergoes pre-training with various pretext tasks on large-scale datasets of image and sentence pairs, resulting in significant performance improvement on downstream tasks including VQA.
We applied our proposed method to these baselines to enable them with the ability of abstention. All the models are maintained with the same experimental settings with the original training protocols.
V-B Experimental Results
Overall Comparison. We benchmark three methods, each with two variants, and show the results in Table V. The observations from this table are three-fold:
@@ V-B Experimental Results
In addition, we also studied the model performance variance with respect to different threshold values on the validation set. The results are demonstrated in Fig. 8. It is evident that UpDn achieves optimal performance with smaller threshold values, such as 10, whereas LXMERT demonstrates improved results with larger threshold values, such as 100. This observation highlights the varying preferences of different models in terms of threshold selection.
Training from Scratch v.s. Fine-tuning In the context of the UNK-VQA dataset, there are generally two training strategies: training a VQA model from scratch and fine-tuning a pre-trained VQA model. To compare their performance, we conducted experiments and presented the results in Fig. 9. The figure demonstrates that for the UpDn model, fine-tuning outperforms end-to-end training by a significant margin. On the other hand, in the case of LXMERT, both training strategies achieve comparable results. Based on these findings, we advocate for the fine-tuning approach when training a VQA model on the UNK-VQA dataset.
Re-evaluation on the Original VQA Dataset In addition to the results on the UNK-VQA dataset, our focus also lies on evaluating the re-trained model’s performance on the original VQA v2 dataset [1]. To explore this, we conducted a re-evaluation of the models trained on UNK-VQA and assessed their performance on VQA v2. The outcomes of this analysis are presented in Table VI. As expected, the model’s performance tends to degrade when evaluated on VQA v2 due to the introduction of new instances from UNK-VQA, which serve as out-of-distribution outliers for VQA v2. Besides this, we have two important findings based on the experiment. 1) Smaller models (UpDn and LXMERT) exhibit a greater degree of performance degradation compared to larger models. This finding suggests that smaller models are more susceptible to perturbations and variations in dataset distribution. 2) The generation model, i.e., BLIP, shows a relatively lower impact after fine-tuning on UNK-VQA. This implies that the generation model is more resilient to the effects of training on UNK-VQA compared to other models. As such, generating the correct answer, rather than simply classifying it, holds the potential to be more advantageous in future approaches.
TABLE VI: Model performance before and after fine-tuning on UNK-VQA. All refers to the overall accuracy, while Y/N, Num., and Other represent sub-categories according to different answer types.

### v4

# UNK-VQA: A Dataset and a Probe into the Abstention Ability of Multi-modal Large Models

@@ I Introduction
Figure 1: Left subfigure: Five perturbation types from our UNK-VQA dataset and their corresponding exemplars. Right subfigure: The unanswerability ratio with respect to each perturbation type.
@@
Despite the crucial function of these benchmarks during the early phase of exploration, recent progress has largely been driven by the emergence of large models [7, 8, 9]. These multi-modal large models overshadow the common training-and-evaluation paradigm in various domains, including VQA. For example, GPT-4 [8] achieves a zero-shot accuracy of $\sim$80on the VQA v2 dataset [1], significantly outperforming most supervised approaches. Pushing the limits on these traditional well-studied benchmarks thus appears to yield marginal contributions. While effective, these large foundation models, however, are notorious for being trustworthy. One demonstration is that they are less capable of abstaining from answering questions that cannot be answered or are beyond their scope of knowledge [10].
@@
This paper contributes a dataset that teaches machines to identify and refuse unanswerable visual questions. Arguably, annotating a large-scale dataset with unanswerable questions is extremely laborious and difficult to control. To address this challenge, we suggest introducing perturbations to existing data and curating hard instances that can potentially deceive models. Based on the most popular VQA v2 dataset [1], we apply five different types of nuanced perturbations to either the given image or the question, as shown in Fig. 1. To assure high quality, we engage the Amazon Mechanical Turk (MTurk) workers to label them manually. In short, the total number of participating workers is $>$4,000, and each instance is annotated by a minimum of three workers. This finally amounts to 10K instances in our dataset, which we name as UNK-VQA. As can be observed from Fig. 1, masking the most relevant object (I-2) often leads to more difficulties in performing question answering. Unlike existing approaches [11, 12, 13] that employ unsupervised learning techniques to address unanswerable visual questions, our proposed dataset enables model training towards more generalizable VQA methods.
@@
We evaluate the performance of multiple multi-modal large models including Otter [14], Open-Flamingo [9], and InstructBLIP [15], as well as a proprietary model GPT-4V. Our findings indicate that: 1) In terms of zero-shot performance, our UNK-VQA dataset reveals limitations in these large models, unlike the impressive results observed in previous general VQA datasets. 2) For few-shot experiments, when we introduce more instances to these models, consistent improvements in their performance can be observed. Additionally, we also propose a simple method to enhance VQA models with the ability of abstention. The key to our method lies in the design of selection functions, which we have developed in several variants, such as a binary classifier and entropy-based methods. Considering the resource-intensive nature of fine-tuning large models, we apply this method to several conventional and recent medium-sized VL Transformer models. Furthermore, we also conduct supervised fine-tuning of the LLaVA model [16] to observe its performance gain on our UNK-VQA.
@@
This dataset reflects the fact that existing multi-modal large models, in comparison to their counterparts in the language domain, are not as omnipotent. It sheds light on various issues beyond the challenge of abstaining from answering unanswerable questions. Problems such as consistency in VQA [17], factuality, and hallucination [8] further emphasize the need for future research in this domain.
@@
  • We introduce a straightforward method for training VQA models that enables them to handle unanswerable questions. Moreover, we show the effectiveness of the proposed method as well as the results of supervised fine-tuning of the LLaVA model.
Figure 2: Illustration of text-based perturbations. We utilize a strong multi-modal model BLIP [18] to generate answers (Base-A) to the perturbed question (Pert-Q). The left example shows that when we replace the anchor noun word with an alternative, the baseline still generates a reasonable answer, even though the modified question becomes unanswerable. Regarding the semantic negation examples, each question can have an uncountable number of potential answers.
@@ II-A Perturbation Procedure and Quality Control
In this work, we introduce five distinct types of perturbations that are applied to either the question or image. We provide a detailed explanation of these perturbations below.
@@
Figure 4: Sunburst distribution of the first four words in the UNK-VQA dataset questions. Most questions begin with the word ‘what’.
@@ II-B Human Labeling
[…]
@@ II-C Data Analysis
Figure 5: Labeling confidence and consensus among three annotators (y-axis: counted numbers). The majority of annotators exhibits a higher level of confidence and can reach a consensus regarding the question answerability.
@@
Figure 6: Left: Interplay between answers (A*) and reasons (R*) for unanswerable questions. The most frequent reasons observed are Being unclear to comprehend, while the most common answer is I don’t know. Right: Answer shift of answerable questions. Notably, many answers shift from the Original ground-truth answers to the Baseline predicted answers.
@@ III Robustness of Large Models towards Unanswerable Questions
Given the remarkable performance of large models across various benchmark datasets, we aim to investigate their effectiveness on our UNK-VQA. Specifically, we focus on examining both publicly available models like OpenFlamingo [9] and a proprietary model, GPT-4V.
III-A Evaluation on Open-sourced Models
III-A1 Evaluation Setup
@@
  • Open-Flamingo-MPT-7B follows similar architecture with Flamingo [9] but incorporates open-sourced vision and language encoder components, specifically CLIP [21] and MPT [28]. The model is trained using a combination of LAION-2B [29] and Multimodal C4 [30], with the weights of both the vision and language encoders frozen.
@@
  • Instruct-BLIP-Vicuna-7B [15] is built on the pre-trained BLIP-2 [7]. It leverages a dataset in an instruction-tuning format that has been adapted from public datasets.
@@
TABLE II: Zero-shot performance of large multi-modal models on three distinct settings.
@@
  • Open-Ended with Hint (OEH) involves an extra hint that indicates the reason for being unanswerable (we collected in Sect. II-B), compared to OE.
@@
Options: A. #opt A B. #opt B C. #opt C D. #opt D
@@
#question If you feel it #unanswerable-reason, you can simply reply “unanswerable”.
Note that in the $k$-shot setting, the above prompts are repeated $k$ times combined with different pre-defined examples, which serve as few-shot prompts.
TABLE III: OoS (Out-of-Scope) Response Ratio of Otte. Generally, we observe that when the model accuracy is higher, the OoS response ratio tends to be lower.
@@
55.05+9.79 55.20+0.41 57.75+3.18
III-A2 Evaluation Results
Zero-shot Performance. We present the zero-shot results in Table II and observe that Instruct-BLIP-Vicuna-13B outperforms other models in both BY and OE settings. Notably, the performance improvements of Instruct-BLIP-Vicuna-13B over Instruct-BLIP-Vicuna-7B highlight the advantages of scaling up the LLM backbone. Another observation is that Open-Flamingo exhibits superior performance compared to the other three models in the MC setting. One possible explanation for this is that, in contrast to BY and OE, MC deviates more from the objective of instruction tuning, e.g., open-domain VQA and image captioning. Consequently, Open-Flamingo receives less training in instruction tuning, which sets it apart from the other models.
@@ III-A2 Evaluation Results
Few-shot Performance. We show the results of this experiment in Fig. 7. One can see that the model performance improves consistently across most experiments. However, one special case with Otter is that the increased number of shots leads to a decline in performance. Upon examining Otter, we found that having more prompted examples often leads to an increase in out-of-scope responses. For example, some models tend to respond I do not know or attempt to explain the unusual nature of the image. These are instances where there is no explicit choice of being answerable or unanswerable in the responses. Additionally, the model’s performance significantly varies based on the number of answerable and unanswerable examples, as illustrated in Table III.
Effectiveness of Post-hint Explanation. Note that in Sec. II-B, we ask each annotator to label the unanswerable reason for each instance. We further investigate whether the post-hint explanations assist in determining the answerability of visual questions. Based on this idea, we provide hints for the unanswerable instances and present the final results in Table IV. The table demonstrates that, in most cases, these post-hints aid the model in comprehending the query more effectively, especially for examples with fewer shots.
@@
III-B Evaluation on Proprietary Model
Initially, we opt to explore the capabilities of the widely used ChatGPT model [^3] https://openai.com/chatgpt.. To adapt the dataset to ChatGPT’s text-only requirement, we employ the strong image caption model BLIP-2 [7] to generate detailed captions for each image. These captions are then used as the supporting document when asking questions to ChatGPT. However, the outcomes are far from convincing. The generated captions failed to accurately depict the image content, often missing crucial objects or neglecting nouns mentioned in the question. Consequently, ChatGPT deemed most questions as ‘unanswerable’, leading to unexpected results. In light of these limitations, we present to directly employ the GPT-4V [^4] https://openai.com/research/gpt-4v-system-card. for our evaluation.
GPT-4V incorporates additional modalities (such as image inputs) into LLMs. Due to budget constraints, we conducted evaluations on a randomly selected subset of 100 instances from our UNK-VQA dataset. These instances encompass all five perturbation types.
Results. The experimental results of GPT-4V are presented in Table V. There are three notable observations from this table: I) GPT-4V exhibits significantly superior performance compared to other open-sourced models in the zero-shot setting. This further proves the superiority of the proprietary GPT-4V over existing open-sourced multi-modal large models. II) GPT-4V benefits little improvement from multiple-shot learning. This aligns with recent findings indicating that current LLMs struggle as few-shot information extractors due to task contamination, particularly for proprietary models [32, 33]. III) There still remains ample room for improvement, underscoring the challenging nature of our UNK-VQA dataset.
TABLE V: GPT-4V results on a subset of UNK-VQA.
Model BY (%) MC (%) OE (%)
0-shot 2-shot 0-shot 2-shot 0-shot 2-shot
GPT-4V 58.0 57.0 38.0 43.0 30.0 28.0
@@ IV-A Preliminary
where $\hat{g}(x)$ indicates the classifier $f$ confidence on the input $x$, and $\theta$ controls the overall prediction versus abstention level.
@@ IV-C Variants of Selective Functions
Classifier-based (CLS) approach involves the introduction of an additional binary classifier, which facilitates the determination of answerability by comparing the predicted score with a predefined threshold $\theta$. Specifically, we re-utilize the fused multi-modal feature $\mathbf{x}$ as inputs to this classifier, represented as $\sigma(\mathbf{W}_{b}\mathbf{x}+\mathbf{b})$, where $\sigma$ represents the sigmoid function.
Entropy-based (ENT) variant compares the entropy of the predicted logits with the threshold $\theta$. If the entropy of the predicted logits, denoted as $H(\mathbf{\pi})=-\sum_{j}^{|\mathcal{A}|}\pi_{j}\log(\pi_{j})$ [^5] In our implementation, we ease the constraint of the infinite range of entropy by comparing the maximum logit with the specified threshold., exceeds $\theta$, indicating a higher degree of randomness, the question is classified as unanswerable. During training, we assign a uniform distribution $\{\frac{1}{|\mathcal{A}|},\frac{1}{|\mathcal{A}|},...\}$ to label the unanswerable questions, thereby reducing the model’s confidence to answer such questions. This approach aims to handle cases where the model should refrain from providing a definitive answer.
TABLE VI: Model performance on UNK-VQA. We did not implement the ENT variant of BLIP because it would have required more computational power due to the generative decoding strategy. Instead, we chose to include an additional label, unanswerable, as a generative candidate during both training and inference.
Model CLS ENT valid test
Accb Acco F1W Accb Acco F1W
@@
V Experiments
We conducted extensive experiments on the newly collected dataset spanning two main aspects: 1) We evaluated the effectiveness of the proposed selective functions; 2) We performed supervised fine-tuning on a strong MMLM - LLaVA [16].
V-A Method Evaluation
V-A1 Experimental Settings
Metrics. For our evaluation, we adopted three metrics:
  • Accb refers to the binary classification accuracy.
  • Acco compares the overlap between the predicted answer and a pre-defined open set with an unanswerable option.
  • Weighted F1 (F1W) is defined as a harmonic mean of the precision Pj and recall Rj ($j$ denotes the $j$-th answer) – $F1^{W}_{j}=2\frac{P_{j}\cdot R_{j}}{P_{j}+P_{j}}$. The final metric takes the answer class imbalance into consideration,
    where $\psi_{j}$ is the ground-truth sample list with answer $j$.
Baselines. We selected three popular baselines to evaluate the effectiveness of the proposed method.
  • LXMERT [38] is built upon the Transformer encoders. It undergoes pre-training with various pretext tasks on large-scale datasets of image and text pairs, resulting in significant improvement on downstream tasks including VQA.
We applied our proposed method to these baselines to enable them with the ability of abstention from unanswerable questions. All the models are maintained with the same experimental settings as the original training protocols.
V-A2 Experimental Results
@@
Overall Comparison. We benchmark three methods, each with two variants, and show the results in Table VI. The observations are three-fold:
@@
In addition, we also studied the model performance variance with respect to different threshold values on the validation set. The results are demonstrated in Fig. 8, which demonstrates that UpDn achieves optimal performance with smaller threshold values, such as 10.
TABLE VII: Model performance before and after fine-tuning on UNK-VQA. All refers to the overall accuracy, while Y/N, Num., and Other represent sub-categories according to different answer types.
@@
Training from Scratch v.s. Fine-tuning. In the context of the UNK-VQA dataset, there are generally two training strategies: training a VQA model from scratch and fine-tuning a pre-trained VQA model. To compare their performance, we conducted experiments and presented the results in Fig. 9. The figure demonstrates that for the UpDn model, fine-tuning outperforms end-to-end training by a significant margin. On the other hand, in the case of LXMERT, both training strategies achieve comparable results. Based on these findings, we advocate for the fine-tuning approach when training a VQA model on the UNK-VQA dataset.
Re-evaluation on the Original Dataset. In addition to the results on UNK-VQA, our focus also lies on evaluating the re-trained model’s performance on the original VQA v2 dataset [1]. To explore this, we conducted a re-evaluation of the models trained on UNK-VQA and assessed their performance on VQA v2. The outcomes of this analysis are presented in Table VII. As expected, the model’s performance tends to degrade when evaluated on VQA v2 due to the introduction of new instances from UNK-VQA, which serve as out-of-distribution outliers for VQA v2. Besides this, we have two important findings based on the experiment. 1) Smaller models (UpDn and LXMERT) exhibit a greater degree of performance degradation compared to larger models. This finding suggests that smaller models are more susceptible to perturbations and variations in dataset distribution. 2) The generation model, i.e., BLIP, shows a relatively lower impact after fine-tuning on UNK-VQA. This implies that the generation model is more resilient to the effects of training on UNK-VQA compared to other models. As such, generating the correct answer, rather than simply classifying it, holds the potential to be more advantageous in future approaches.
V-B Supervised Fine-tuning of LLaVA
LLaVA [16] represents a novel end-to-end trained large multi-modal model that combines a vision encoder and an LLM for general-purpose visual and language understanding. It connects the visual patch-level features from CLIP [21] and LLM with a simple linear layer, and is fine-tuned on a multi-modal instruction tuning dataset. In our experiment, we utilized a more advanced LLaVA-1.6 model and performed supervised fine-tuning on our UNK-VQA dataset. Specifically, we appended a prompt to each question - Answer the question using a single word or phrase. If you feel you cannot answer this question, simply reply “Unanswerable”. Regarding the LLM, we selected both a weaker Vicuna-13B model and a recent strong Mistral-7B model.
Experimental Results. From the results in Table VIII, we can observe that: I) The LLaVA models exhibit a significant performance advantage over traditional models, as demonstrated in Table VI, owing to their pre-training on more extensive datasets. II) Both LLaVA models, utilizing different LLMs, show substantial improvement through supervised fine-tuning. III) Despite its smaller model size, the stronger Mistral-7B LLM notably outperforms the Vicuna-13B model. This highlights the pivotal role of LLM capability in advancing multi-modal large models in UNK-VQA.
TABLE VIII: Supervised fine-tuning results of LLaVA on UNK-VQA.
Model LLM FT BY (%) OE (%)
LLaVA-1.6 [16] Vicuna-13B [31] 48.18 21.53
58.00 52.49
Mistral-7B [39] 56.98 49.30
56.35 54.16